# SSL-OTA: Unveiling Backdoor Threats in Self-Supervised Learning for Object Detection

### v1

@@ 4.1 Experiment settings
Trigger. We employed publicly released HTBA triggers [20], which are square triggers generated by bilinear interpolation, adjusting random 4×4 RGB images to the required patch size. The HTBA triggers are indexed from 10 to 19, and we utilized the trigger with index 10.
@@
Settings. In our experiments, we utilized a ResNet-50 architecture for the pre-trained encoder, opting for the results of 200 rounds of pre-training with MoCo-v1 on the ImageNet dataset [21]. The detector models, Faster R-CNN [16], and Mask R-CNN [22], utilized the R50-C4 backbone [22], with BN tuning following the guidelines in [23]. The R50-C4 backbone is similar to those available in [23], concluding its main network at the conv4 stage, while the box prediction head comprises the conv5 stage (including global pooling) and BN layers.
@@
During downstream fine-tuning, we loaded the pre-trained encoder’s weights as input for the partial detector backbone network, subjecting all layers to end-to-end fine-tuning. The image scale ranged from [480, 800] pixels during training, set to 800 pixels during inference. We employed an optimizer strategy consisting of two stages, initializing the learning rate at 0.02 and adjusting it at iterations 18K and 22K, with a maximum of 24K iterations. All settings remained consistent.
@@
We fine-tune on the VOC train2007 or a portion of the MSCOCO train2017($\sim$58k images) at a 0.5% poisoning rate in the targeted attack classes. In the case of DSBA, we employed a portion of the MSCOCO train2017($\sim$60k images) to train a clean object detector. In the backdoor injection phase, we used this dataset to construct a shadow dataset. Specifically, we added an HTBA trigger of size 29*29 to the center of the ”person” class’s bbox and modified its annotation category to ”dog.”
@@
Evaluation Metrics. Inspired by research on backdoor attacks in supervised learning for object detection [11, 12], we introduced effective evaluation metrics in the SSL context to quantify the impact of backdoor attacks on the performance of object detection. It is important to note that we utilize detection metrics, namely AP and mAP at IoU = 0.5, denoted as $AP_{50}$ and $mAP_{50}$.
@@
To ensure that the behavior of the $M_{infected}$ on benign inputs across all settings is similar to the clean model $M_{benign}$, we use the mAP obtained on the clean test dataset $D_{test,benign}$ as $mAP_{benign}$. Additionally, we obtain the AP for the target class $t$ on $D_{test,benign}$ as $AP_{benin}$. To validate that $M_{infected}$ successfully predicts bboxes as the target class, we calculate $AP_{attack}$ for the target class $t$ on the backdoored test dataset $D_{test,backdoored}$. A high $AP_{attack}$ indicates that, due to the presence of triggers, more bboxes are predicted as the target class with high confidence scores.
@@
Furthermore, we construct a hybrid dataset for backdoor evaluation, denoted as $D_{\text{test, backdoored }+\text{ benign }}=\left\{\left(x_{\text{backdoored }},y\right)\right\}$, combining backdoored images $x_{backdoored}$ from $D_{test,backdoored}$ with ground-truth labels $y$ from $D_{test,benign}$. To demonstrate that boxes are changed to the target class, we compute $AP_{attack+benign}$ for the target class $t$ in $D_{test,backdoored+benign}$. Changes to the target category for bboxes are considered false positives, marked against the ground-truth labels $y$, resulting in a low $AP_{attack+benign}$.
@@
To verify that $M_{infected}$ does not misclassify bboxes with non-target classes, we performed calculations of $mAP_{part\_attack+benign}$ on the $D_{test,part\_backdoored+benign}$. This dataset consists of five randomly selected categories, each with an added trigger pattern.We exclude the scenario where trigger patterns are added to all categories, as many smaller targets in the VOC2007 and MSCOCO datasets would be entirely obscured by trigger patterns. Regarding OTA, only bboxes belonging to the attacked category, misclassified as target classes, do not significantly impact $mAP_{part\_attack+benign}$, given the presence of numerous other categories.
4.2 Attack Performance
We evaluated $M_{infected}$ using various metrics mentioned in Section4.1. For both NA and DSBA, we employed the same settings: a poisoning rate ($P$) of 0.5%, a trigger size of 29*29, a trigger ratio of 1, and ”person” as the target category. We present the results of the two attacks on the shadow dataset with different or similar distributions to the downstream dataset in Table 1. Additionally, we displayed the evaluation results of $M_{benign}$ in Table 2 for comparison with $M_{infected}$.
On the Fast R-CNN model, both NA and DSBA demonstrated strong attack capabilities on the VOC2007 dataset. Notably, DSBA achieved an ASR of 86.55% on this dataset, significantly higher than NA. As an effective and covert backdoor implantation method in SSL contexts, DSBA exhibited a higher ASR and comparable benign model utility impact. On the MSCOCO dataset, both attack methods showed high ASRs, indicating that NA, even with a very low poisoning rate (e.g., 0.5%), remains effective in complex datasets. In comparison, DSBA’s Attack Precision ($AP_{attack}$) and ASR improved on MSCOCO, further emphasizing its attack efficacy in complex scenarios.
In most cases, the overall utility loss of both attacks, when comparing $M_{infected}$ to $M_{benign}$, was less than 1%. DSBA caused a slight decrease in the $mAP_{benign}$ on the $\langle MaskR-CNN,MSCOCO\rangle$. However, DSBA did not exhibit the expected significant advantage on the more complex MSCOCO dataset. This might be due to DSBA’s involvement with pre-trained encoders and its complex dual-attack strategy, which is more effective in simpler datasets like VOC2007. However, this advantage might be mitigated in MSCOCO, which contains numerous small objects and detailed annotations, due to the inherent noise and diversity of the data. Overall, both NA and DSBA are capable of efficient attacks across different datasets and model architectures with minimal impact on model performance. This highlights the potential risks and stealthiness of such attacks in real-world scenarios.
Table 2: Performance of benign model $M_{benign}$ for $mAP_{benign}$/$AP_{benign}$ comparison with infected model $M_{infected}$.
@@ 4.2 Attack Performance
4.3 Ablation Study
To investigate the different components of the backdoor attacks we introduced, we analyzed the influence of poisoning rate ($P$) and trigger size $(W_{t},H_{t})$ on both two attacks. Apart from the parameters under study, all other settings remain consistent with those in Section 4.1. We report both the attack and utility performance in Figure 2.
@@ 4.3 Ablation Study
Fig. 2: The effects of poisoning rates and trigger size. (a)The impact of poisoning rate on NA. (b)The impact of poisoning rate on DSBA. (c)The impact of trigger size on NA . (d)The impact of trigger size on DSBA.
As shown in Figure 2 and Figure 2, the DSBA exhibits significant effectiveness (ASR and $AP_{attack}$) at lower poisoning rates, whereas the NA requires a higher poisoning rate to achieve a comparable ASR. Moreover, as depicted in Figure 2 and Figure 2, DSBA demonstrates low sensitivity to trigger size, maintaining a stable ASR across changes in trigger dimensions, and achieves a high ASR (87.74%) with a trigger size of 10*10. In contrast, the ASR of NA increases with the size of the trigger, with smaller dimensions (such as 5*5) resulting in a significant drop in attack performance. This suggests that DSBA offers greater flexibility and robustness in executing effective backdoor attacks. Additionally, both attacks have a negligible impact on the model’s utility (as indicated by metrics such as $mAP_{benign}$), suggesting their ability to covertly reside within the model without detection. Overall, while NA is effective under certain conditions, DSBA provides superior performance and greater adaptability across various aspects.
5 Conclusion
@@
In this work, we investigate, for the first time, backdoor attacks on downstream tasks of object detection in the SSL scenario, elucidating the potential backdoor threats faced by object detection in this context. In the OTA, we propose a simple yet effective backdoor attack strategy (NA) and further introduce an enhanced attack strategy (DSBA). We define appropriate metrics for evaluating attack performance, demonstrating the effectiveness and utility of both attack strategies. Additionally, we delve into an in-depth exploration of the impact of various parameters and triggering factors on the efficacy of the attacks. Our approach serves as a valuable tool for assessing the backdoor robustness of object detectors, contributing to the design of more secure models.

### v2

@@ IV-A Experiment settings
Trigger. We employ publicly released HTBA triggers [23], which are square triggers generated by bilinear interpolation, adjusting random 4×4 RGB images to the required patch size. The HTBA triggers are indexed from 10 to 19, and we utilize the trigger with index 10.
@@
Settings. In our experiments, we utilize a ResNet-50 architecture for the pre-trained encoder, leveraging the results of 200 rounds of pre-training with MoCo-v1 on the ImageNet dataset [24]. The detector models, Faster R-CNN [17], and Mask R-CNN [25], employ the R50-C4 backbone [25], with batch normalization (BN) tuning as outlined in  [26]. The R50-C4 backbone is similar to those available in [26], concluding its main network at the conv4 stage, while the box prediction head comprises the conv5 stage (including global pooling) and additional BN layers.
@@
During downstream fine-tuning, we load the pre-trained encoder’s weights as input for the partial detector backbone network, subjecting all layers to end-to-end fine-tuning. The image scale ranged from [480, 800] pixels during training, set to 800 pixels during inference. We employ an optimizer strategy consisting of two stages, initializing the learning rate at 0.02 and adjusting it at iterations 18K and 22K, with a maximum of 24K iterations. All settings remain consistent.
@@
We fine-tune on the VOC train2007 or a portion of the MSCOCO train2017 ($\sim$58k images) at a 0.5% poisoning rate in the targeted attack classes. In the case of DSBA, we employ a portion of the MSCOCO train2017 ($\sim$60k images) to train a clean object detector. In the backdoor injection phase, we use this dataset to construct a shadow dataset. Specifically, we add an HTBA trigger of size 29*29 to the center of the “person” class’s bbox and modify its annotation category to “dog”.
@@
Evaluation Metrics. Inspired by research on backdoor attacks in supervised learning for object detection [12, 13], we introduce effective evaluation metrics in the SSL context to quantify the impact of backdoor attacks on the performance of object detection. It is important to note that we utilize detection metrics, namely AP and mAP at IoU = 0.5, denoted as $AP_{50}$ and $mAP_{50}$.
@@
To ensure that the behavior of the $M_{infected}$ on benign inputs across all settings is similar to the clean model $M_{benign}$, we use the mAP obtained on the clean test dataset $D_{test,benign}$ as $mAP_{benign}$. Additionally, we obtain the AP for the target class $c_{t}$ on $D_{test,benign}$ as $AP_{benign}$. To validate that $M_{infected}$ successfully predicts bboxes as the target class, we calculate $AP_{attack}$ for the target class $c_{t}$ on the backdoored test dataset $D_{test,backdoored}$. A high $AP_{attack}$ indicates that, due to the presence of triggers, more bboxes are predicted as the target class with high confidence scores.
@@
Furthermore, we construct a hybrid dataset for backdoor evaluation, denoted as $D_{test,backdoored+benign}=\left\{\left(\boldsymbol{x}_{backdoored},\boldsymbol{y}\right)\right\}$, combining backdoored images $\boldsymbol{x}_{backdoored}$ from $D_{test,backdoored}$ with ground-truth labels $\boldsymbol{y}$ from $D_{test,benign}$. To convey that boxes are changed to the target class, we compute $AP_{attack+benign}$ for the target class $c_{t}$ in $D_{test,backdoored+benign}$. Changes to the target category for bboxes are considered false positives, marked against the ground-truth labels $\boldsymbol{y}$, resulting in a low $AP_{attack+benign}$.
@@
To verify that $M_{infected}$ does not misclassify bboxes with non-target classes, we calculate $mAP_{part\_attack+benign}$ on the $D_{test,part\_backdoored+benign}$. This test dataset consists of five randomly selected categories, each with an added trigger pattern. We exclude the scenario where trigger patterns are added to all categories, as many smaller targets in the VOC2007 and MSCOCO datasets would be entirely obscured by trigger patterns. Regarding SSL-OTA, only bboxes belonging to the attacked category, are misclassified as target classes, and do not significantly impact $mAP_{part\_attack+benign}$, given the presence of numerous other categories.
TABLE II: Performance (%) Comparison Analysis of $M_{benign}$ and $M_{infected}$.
@@
$ASR$ 0.003 0.001 0.001
@@
Figure 3: Impact of poisoning rate, trigger size, and trigger rate on attack effectiveness. (a) The impact of poisoning rate on NA. (b) The impact of trigger size on NA. (c) The impact of trigger ratio on NA. (d) The impact of poisoning rate on DSBA. (e) The impact of trigger size on DSBA. (f) The impact of trigger ratio on DSBA.
IV-B Attack Performance
We evaluate $M_{infected}$ using various metrics mentioned in SectionIV-A. For both NA and DSBA, we employ the same settings: a notably low poisoning rate ($P$) of 0.5%, a trigger size of 29*29, a trigger ratio of 1, and “person” as the target category. We present the detailed results of the two attacks on the shadow dataset with different or similar distributions to the downstream dataset in Table I. Additionally, we display the evaluation results of $M_{benign}$ in Table II for comparison with $M_{infected}$.
On the Fast R-CNN model, both NA and DSBA demonstrated strong attack capabilities on the VOC2007 dataset. Notably, DSBA achieved an ASR of 86.55% on this dataset, significantly higher than NA. As an effective and covert backdoor implantation method in SSL contexts, DSBA exhibits a higher ASR and comparable benign model utility impact. On the MSCOCO dataset, both attacks show high ASR, indicating that NA, even with a very low poisoning rate (e.g., 0.5%), remains effective in complex datasets. In comparison, DSBA’s Attack Precision ($AP_{attack}$) and ASR improved on MSCOCO, further emphasizing its attack efficacy in complex scenarios.
[…]
IV-C Ablation Study
To investigate the different components of the backdoor attacks we introduced, we analyze the influence of poisoning rate ($P$), trigger size $(W_{t},H_{t})$, trigger ratio $\alpha$, target class $c_{t}$, and trigger position on both two attacks. Apart from the parameters under study, all other settings remain consistent with those in Section IV-A. Each ablation study altered a single parameter to observe its effects, and we report both the attack and utility performance in Figure 3.
The effects of poisoning rates. As shown in Figure 3 and Figure 3, we explored the impact of poisoning ratios in the downstream fine-tuned dataset on the effectiveness of attacks. We observed that the attack performance of NA, as indicated by metrics such as ASR score and $AP_{attack}$, increases with the rise in poisoning ratio $P$, showing significant overall changes. In contrast, the attack performance of DSBA is relatively less sensitive to changes in the poisoning ratio. Even at lower poisoning levels, DSBA demonstrates substantial effectiveness. For instance, at a poisoning ratio of merely 0.001, DSBA’s $AP_{attack}$ reaches 86.3%, and its ASR attains 74.8%. Conversely, NA requires a higher poisoning rate to achieve a similar level of ASR. Regarding the model’s practical performance, we noted that parameters such as $mAP_{benign}$ slightly decrease as the poisoning ratio increases, with the reduction in NA being slightly greater than that in DSBA. This indicates that both attacks can maintain high attack performance while ensuring the model’s utility.
The effects of trigger size. We also demonstrated the impact of trigger size on attack performance, as shown in Figures 3 and Figure 3. We found that larger trigger sizes enhance attack effectiveness. The success rate of NA increases with the size of the trigger. A small trigger size (such as $5*5$) significantly reduces attack performance, whereas a size of $10*10$ ensures relatively stable attack outcomes. In contrast, DSBA exhibits low sensitivity to trigger size, maintains a steady ASR across different trigger sizes, and achieves a high ASR of 85.48% with a trigger size of $10*10$. Notably, DSBA maintains commendable attack performance even with a minimal trigger size of $5*5$, achieving an ASR of 67.83% and an $mAP_{benign}$ of 71.18%. This contributes to making the trigger more difficult to detect, thereby enhancing the stealth of the attack.
The effects of trigger ratio. Furthermore, our study investigated the impact of trigger rates $\alpha$ on attack outcomes, as illustrated in Figures 3 and 3. We discovered that higher trigger rates, denoted by $\alpha$, have a minimal effect on metrics such as ASR and $mAP_{benign}$ across both attacks. In most scenarios, regardless of the trigger ratio settings, the ASR for NA remains close to 73%, while for DSBA, it hovers around 85%, with both maintaining a $mAP_{benign}$ similar to that of a clean model (approximately 70%). This indicates that attackers could employ minimal trigger ratios (e.g., $\alpha=0.1$) to render the trigger virtually invisible in images, suggesting a high potential for stealth in attack implementation.
The effects of target class and trigger position. To assess the impact of target class variation, we modified the target class from “dog” to “sheep,” a category with fewer instances. According to the results presented in Table III(a), the performance of the NA and DSBA methods is not adversely affected by the reduced number of target class objects. Furthermore, to demonstrate that the placement of the trigger does not influence the outcome of the attack, we altered the position of the trigger to random locations within the images of the poisoned dataset, rather than centering it within the bbox. The findings, as shown in Table III(b), indicate that this alteration does not affect the efficacy of the attack, paralleling the results observed in Table I.
[…]
(a) Target class t = “sheep” class.
Attack Method NA DSBA
$mAP_{benign}$ 69.62 71.14
$AP_{benign}$ 50.98 48.61
$mAP_{attack}$uparrow 63.89 68.20
$AP_{attack}$uparrow 76.65 89.53
$mAP_{attack+benign}$ 61.62 64.46
$AP_{attack+benign}$downarrow 0.30 0.23
$mAP_{part\_attack+benign}$ 60.08 62.83
$ASR$uparrow 74.52 87.04
(b) Random triggers’ locations.
Attack Method NA DSBA
$mAP_{benign}$ 69.97 72.44
$AP_{benign}$ 69.49 72.25
$mAP_{attack}$uparrow 61.52 65.88
$AP_{attack}$uparrow 79.74 88.05
$mAP_{attack+benign}$ 62.52 64.92
$AP_{attack+benign}$downarrow 1.67 1.74
$mAP_{part\_attack+benign}$ 60.34 62.18
$ASR$uparrow 71.44 82.12
Figure 4: The resistance to model pruning. (a) The effect of NA against model pruning. (b) The effect of DSBA against model pruning.
Overall, DSBA demonstrates greater flexibility and robustness in executing effective backdoor attacks, offering superior performance and versatility across various aspects compared to NA, which is effective under specific conditions. Moreover, both attacks minimally impact model utility (as indicated by metrics such as $mAP_{benign}$), suggesting they can successfully remain undetected within the model.
IV-D The Resistance to Model Pruning
Although many defense or detection methods have achieved significant success in addressing backdoor attacks on image classification, these methods face numerous challenges when applied to object detection tasks. First, methods [27, 28] that predict the distribution of backdoor triggers through generative modeling or neuron reverse engineering typically assume that the model is a simple neural network. In contrast, object detection models consist of multiple complex components, making it difficult to apply these methods directly. Furthermore, the output of image classification models is a single predicted category, while object detection models output multiple objects along with their positions and categories. This diversity in output further complicates the detection of backdoor triggers.
In the SSL scenario, there are currently no specific defense measures against backdoor attacks on object detection models. Therefore, we consider using general backdoor defense strategies to evaluate the robustness of our proposed attack. Model pruning [29, 30] is a typical backdoor defense method that can be widely applied to various tasks. It reduces the impact of backdoor triggers by removing neurons with low activation rates in benign datasets. In this section, we will evaluate whether our attack can withstand this potential defense. Specifically, after model training is complete, we perform model pruning during the testing phase. Using 40% of the benign test samples, we prune neurons with the lowest activation values by comparing the model’s performance on the backdoor test dataset and the clean dataset. As shown in Figure 4, as the pruning rate increases, the $AP_{attack}$ values for both attacks decrease, but the $mAP$ values for benign samples also decrease rather than increase. This indicates that model pruning has limited effectiveness in object detection tasks. Object detection models need to retain more neurons to handle complex detection tasks, and pruning can lead to a significant decline in model performance (e.g., $mAP$). In summary, our method can withstand model pruning.
V Related Work
V-A Backdoor Attacks
Backdoor attacks are an emerging topic with potential development prospects in the field of network security, incorporating specially designed triggers into inputs. This compels the model embedded with a backdoor to classify it into a predefined target category with high confidence. Importantly, the backdoor model can still function normally and produce almost the same accuracy as an uninfected model even after the removal of the trigger from the inputs.
In the field of supervised learning, backdoor attacks targeting image classifiers have been broadly explored by researchers. These attacks are designed to misclassify images containing triggers chosen by the attacker, typically by contaminating the training dataset. Such forms of attacks were initially studied in the context of supervised learning [31, 32]. More advanced and difficult-to-detect methods of attack have been proposed in [33, 23, 34, 35], exploring covert attacks under various scenarios. For instance, Quiring et al. [33] investigated attacks targeting image scaling, while Saha et al. [23] assumed that attackers are aware of the model structure. In [34], researchers utilized image structures to create covert trigger areas and embedded them into images using deep injection networks, achieving covert attacks. Similarly, a method involving the conversion from RGB to YUV channels as well as DCT transformations for embedding triggers was proposed in [35].
The area of SSL offers a unique perspective for backdoor attacks. Recent studies, such as those by Saha et al. [8], have demonstrated the embedding of backdoors into the unlabeled datasets of SSL without damaging the training phases of image encoders and downstream classifiers. Building upon this, Li et al. [9] defined trigger patterns as specific disturbances in the frequency domain. However, these attacks have shown shortcomings in terms of attack efficacy and model performance. Other concurrent works [36, 7] have attempted to improve the ASR. Specifically, BadEncoder [7] focuses on executing precise backdoor attacks in SSL encoders by compromising encoder training, assuming access to a clean, pre-trained encoder. This approach has been proven to achieve more than 98% ASR on the ImageNet-100 dataset. Research into backdoor attacks and their defense/detection methods has been extensively explored across multiple fields, including image recognition, video recognition, natural language processing (e.g., sentiment classification, toxicity detection, spam detection), and federated learning. However, existing backdoor attacks on self-supervised pre-trained models have largely focused on image classification tasks, with such attacks on object detection tasks yet to be explored.
V-B Object Detection Using SSL
Significant progress has been made in the field of SSL in recent years. However, most methods [1, 19] are primarily designed for image classification, though they also demonstrate transferability to object detection tasks. For instance, the representations learned by MoCo [1] can be effectively transferred to downstream tasks. Compared to supervised pretraining on ImageNet, MoCo has shown superior transfer capabilities in various object detection tasks.
Although direct applications of SSL in object detection have been less common, recent works like DenseCL [37], InsLoc [38], and PatchReID [39] have begun to explore SSL pretraining tasks specifically designed for object detection. DenseCL enhances local feature learning by conducting dense contrastive learning at the pixel level. InsLoc combines instance-level contrastive learning, resulting in superior instance-level feature representations. PatchReID employs a local region contrast strategy, further improving the model’s performance in fine-grained feature learning for object detection. These methods leverage contrastive learning to improve object detection performance, indicating that SSL techniques are gradually covering a broader range of tasks within the computer vision field.
VI Conclusion
@@
In this work, we investigate, for the first time, backdoor attacks on downstream tasks of object detection in the SSL scenario, elucidating the potential backdoor threats faced by object detection in this context. In the SSL-OTA, we propose a simple yet effective backdoor attack strategy (NA) and further introduce an enhanced attack strategy (DSBA). We define appropriate metrics for evaluating attack performance, demonstrating the effectiveness and utility of both attack strategies. Additionally, we delve into an in-depth exploration of the impact of various parameters and triggering factors on the efficacy of the attacks. Our approach serves as a valuable tool for assessing the backdoor robustness of object detectors, contributing to the design of more secure models.